# Diversifying by Intent in Recommender Systems

### v1

@@ 4. Experiments
We conduct experiments on one of the world’s largest content recommendation platforms, serving billions of users everyday. Similar to many industrial recommendation platforms, the recommender system adopted by the platform for the landing page consists of three stages, including candidate generation, multi-tasking ranking, and whole-page optimization (Yi et al., 2023; Ma et al., 2018; Wilhelm et al., 2018). Our intent diversification framework is applied as the last step in the whole-page optimization stage, taking the scores from the original multi-stage system as $s_{ij}$. Using live A/B testing, we compare the performance of the production recommender system with (treatment) and without (control) our proposed intent diversification framework.
@@ 4.1. Live Experiment Results
Figure 3 summarizes the live experiment metrics. The differences are reported as percentage changes on the metrics of the treatment group with respect to the control. Due to business compliance reasons, the absolute values of the metrics and the x-axis (representing the time horizon of the experiment) have been omitted. The top-line business metric, overall user enjoyment, is a metric combining multiple aspects of user experience in engagement, satisfaction and content quality, and shows a significant improvement of 0.09% (Fig.3(a)) under the intent diversification treatment. At the same time, Daily Active Users (DAU) is improved by 0.05% which is also statistically significant (Fig.3(b)). This is accompanied by the upward trend in the improvement in consumption on the landing page (Fig.3(c)), indicating that the users are increasingly engaging with the platform. Furthermore, there is also an upward trend in satisfied users (Fig.3(d)), suggesting that users are not only increasing their engagement with the platform but also expressing higher satisfaction levels. The improvement in DAU suggests that users are returning to the platform more frequently, indicative of an improved long-term user experience.
@@ 4.2. Understanding the Improvements
To delve deeper into the sources of gain from the intent diversification framework, we conducted additional analyses to examine how users’ exploration behavior evolves over time and its impact on the diversity of their consumption patterns. Recall that the user exploration intent is defined as the user consuming contents from a provider that they have not engaged with before. By explicitly incorporating it into diversification, we see a significant improvement in user engagement with novel providers (Fig.4(a)). To check whether these engagements are high-quality engagements, the platform also measures “repeated exploration”, which is defined as the user visiting the same novel provider multiple times. This can be viewed as a measure for high-quality exploration behavior. Figure 4(b) shows that there is also a significant improvement in repeated exploration, indicating an improvement in users finding high-quality exploratory contents. To measure the diversity of users’ consumption patterns, we also examined the unique topic clusters that users interacted with [^4] These topic clusters, derived from an unsupervised clustering algorithm based on observed co-occurrence patterns on the platform (Chen et al., 2021), indicate semantic similarities between items within the same cluster.. Notably, Fig.4(c) demonstrates a substantial improvement in the average number of unique topic clusters consumed per user. This finding suggests that the intent diversification strategy, by encouraging users to engage with a wider variety of novel content, leads to increased consumption diversity.
@@
(a) Novel content provider engagement.
@@
Our proposed intent diversification framework is a personalized treatment based on a personalized prediction of intent propensities for each user. To understand the effect of personalization, we conducted further analysis by comparing the treatment (with intent diversification) and control (without intent diversification) for users with different levels of exploration intent $p(Exploration)$. In particular, we looked at the following metrics: (1) Novel impressions: Number of impressions on the landing page that are from novel content providers, with whom the user has not previously engaged [^5] Note that the novelty of content providers is a personalized notion. A novel content provider for user A may not be a novel content provider for user B.. This measures the overall novelty of the recommendations. (2) Novel consumptions: Number of consumed items on the landing page that are from novel content providers. (3) Novel click-through rate (CTR): Click-through-rates of the novel recommended items. This measures the relevance of the novel recommendations. Figure 5(a) - 5(c) shows the differences between treatment and control in these metrics. We observe that for pages with higher exploration intent, the treatment (our intent diversification framework) is recommending more novel items (Fig.5(a)) than control, and the users are more likely to engaging with those novel contents (Fig.5(b) and Fig.5(c)).
All three metrics in Fig.5 exhibit a consistent monotonic trend. This suggests that our intent diversification framework improves personalization with respect to users’ exploration intent. Specifically, for pages with a higher propensity towards exploration, our framework not only recommends an increasing proportion of high-quality novel items but also observes increased user engagement with these recommendations, as evidenced by the difference in CTR (Fig.5(c)). For example, at about 80% percentile of $p(Exploration)$, our proposed framework is showing about the same amount of novel recommendations (Fig.5(a)), but results in 4% increase in consumptions from novel contents (Fig.5(b)). This suggest that our framework is more efficient in recommending contents that are both novel and relevant to the users who are in exploration mode. Conversely, for users with lower exploration intent, the recommendation of such novel items decreases and the users are less likely to engage with them. In other words, our intent diversification framework is able to further personalize the recommendation results to align them with varying levels of user exploration intents.
@@
4.3. Understanding User Exploration Intent
The user exploration intent model has an AUC of 0.73 and is well calibrated with an average prediction to label ratio of 0.97 (Fig.6). To understand what signals are captured by the user exploration intent model, we conducted analysis to study how each feature is correlated with the $p(Exploration)$ prediction and rank by the order of the absolute value of the correlations. Table 1 summarizes the features that have the highest absolute correlations with the predicted user exploration intent. We see that features such as “current session length” and “number of consumptions in the current session” are positively correlated with $p(Exploration)$. This is interesting and expected, as it suggests that users are more likely to explore when they have caught up existing interests (e.g. finishing consuming the newest episode of a drama series) in the current session. Features such as “average completion ratio of past consumed items” and “average total length of past consumed items” are negatively correlated with $p(Exploration)$, indicating that users who are less patient with each consumption (i.e. more likely to consume shorter items or quit before finishing the consumption) are more likely to explore. These findings are also aligned with the qualitative insights from user interviews conducted at the company.
@@ 4.3. Understanding User Exploration Intent
Number of content providers interacted with
@@
Table 1. Feature analysis of user exploration intent model.
@@ 4.4. Effect of the Strength of Intent Diversification
Although diversity is a desired property for recommendation results, it has been established that diversity usually hurts the overall relevance of the set of recommended items (Chen et al., 2021). Therefore, the goal of the recommender system is to achieve the optimal trade-off between diversity and relevance in order to optimize long-term user experience. In our intent diversification framework, the strength of diversification is controlled by the hyperparameter $\gamma$ in Eq.(7). To understand the effect of the different strengths of intent diversification, we conduct additional experiments with varying values of $\gamma$, and measure the following four metrics: (1) Diversity, measured by number of unique clusters consumed per user; (2) Novelty, measured by novel content provider engagement; (3) Relevance, which is the average of the relevance score (predicted by a separate ML model) of all recommended items on the landing page; (4) Long-Term Metric (DAU), which is the Daily Active User (DAU) metric.
@@ 5. Conclusion
In this work, we propose a framework to diversify recommendation results based on user intents. By adapting an algorithm originally designed for search results diversification, we propose to explicitly incorporate the user’s propensities to different intents in the final stage of a recommender system, to ensure that different intents are represented in the recommendations. Through live experiments on one of the world’s largest content recommendation platforms, we demonstrate the effectiveness of our proposed framework in assisting users towards their intents and improving their long-term experience on the platform. In particular, when incorporating user exploration intent, we see a significant improvement in long-term metrics such as overall user enjoyment and Daily Active Users. At the same time, we also see an increasing improvement in exploration-related metrics over time, which showcase the connection between user exploration and long-term user experience.
@@
Our proposed intent diversification framework is generalizable and readily applicable to any existing recommender systems. It works with any number of intents and can be applied either as a standalone diversification layer or combined with existing whole-page optimization algorithms. In the live experiments, we selected user exploration intent based on the insights on the strong connection between exploration and long-term user experience from past work (Chen et al., 2021). A future research direction is to automatically learn a taxonomy of user intents in the recommendation context that can directly serve as input to the intent diversification framework toward optimizing long-term user experience on the platform.

### v3

@@ 4. Experiments
4.1. Experiment Setup
We conduct experiments on YouTube, the world’s largest video recommendation platform, serving billions of users everyday. Similar to many industrial recommendation platforms, the recommender system adopted by the platform for the landing page consists of three stages, including candidate generation, multi-tasking ranking, and whole-page optimization (Yi et al., 2023; Ma et al., 2018; Wilhelm et al., 2018). Our intent diversification framework is applied as the last step in the whole-page optimization stage, taking the scores from the original multi-stage system as $s_{ij}$. Using live A/B testing, we compare the performance of the production recommender system with (treatment) and without (control) our proposed intent diversification framework. The experiments were run for a period considered long term by the company [^4] Unfortunately, we cannot disclose the exact time period due to business compliance reasons., allowing us to observe the effects on long-term user experience.
While our proposed intent diversification framework works with any number or types of intents, in the first set of live experiments we selected two intents that are critical to the platform: exploration intent and familiarity intent. Following the intent label definition in Section 3.2, a user is considered to have an exploration intent if they prefer to consume content from an unseen creator (i.e., watch a video from a channel they have not seen before), and a familiarity intent if they prefer content from familiar creators (i.e., watch a video from channels they have seen before). The reason we selected these two intents is that they are closely connected with user exploration in recommender systems, which is defined as identifying unknown user interests or introducing users to new interests (Herlocker et al., 2004). Chen et al. (2021) showed that user exploration is tightly connected to long-term user experience on content recommendation platforms. In addition, incorporating a user’s exploration intent in the diversification stage allows the system to dynamically adjust the diversity level based on whether the user seeks explorational content. This approach enables the diversification stage to adapt to user preferences in real time, effectively implementing a form of “learning to diversify” in recommender systems. Motivated by these considerations, we chose user exploration and familiarity intents for our intent diversification framework to optimize long-term user experience. The intent model is retrained every 6 hours with the latest observations to capture any dynamics in the intent space.
In Section 4.2 below, we present live A/B testing results on these intents. In Appendix A.1, we present live A/B testing results on a different set of intents, specifically creator-level intents and visit length intents, to showcase the generalizability and scalability of our proposed framework to a large number of intents.
@@ 4.2. Live Experiment Results
Figure 3 summarizes the live experiment metrics with the selected exploration and familiarity intents. The differences are reported as percentage changes on the metrics of the treatment group with respect to the control. Due to business compliance reasons, the absolute values of the metrics and the x-axis (representing the time horizon of the experiment) have been omitted. The top-line business metric, overall user enjoyment [^5] This is an objective and well-defined measure on high-quality user engagement on the platform. It combines feedback based on time spent, survey, and content quality guardrail etc. We have kept the name of the metric vague due to business compliance reasons., shows a significant improvement of 0.09% (Fig.3(a)) under the intent diversification treatment. At the same time, Daily Active Users (DAU) is improved by 0.05% which is also statistically significant (Fig.3(b)). This is accompanied by the upward trend in the improvement in consumption on the landing page (Fig.3(c)), indicating that the users are increasingly engaging with the platform. Furthermore, there is also an upward trend in satisfied users (Fig.3(d)), suggesting that users are not only increasing their engagement with the platform but also expressing higher satisfaction levels. The improvement in DAU suggests that users are returning to the platform more frequently, implying an improved long-term user experience.
@@ 4.3. Understanding the Improvements
To delve deeper into the sources of gain from the intent diversification framework, we conducted additional analyses to examine how users’ exploration behavior evolves over time and its impact on the diversity of their consumption patterns. In addition to the improvements in top business metrics in Section 4.2, we also see a significant improvement in user engagement with novel creators (Fig.4(a)). To check whether these engagements are high-quality engagements, the platform also measures “repeated exploration”, which is defined as the user visiting the same novel creator multiple times. This can be viewed as a measure for high-quality exploration behavior. Figure 4(b) shows that there is also a significant improvement in repeated exploration, indicating an improvement in users finding high-quality exploratory contents. To measure the diversity of users’ consumption patterns, we also examined the unique topic clusters that users interacted with [^6] These topic clusters, derived from an unsupervised clustering algorithm based on observed co-occurrence patterns on the platform (Chen et al., 2021), indicate semantic similarities between items within the same cluster.. Notably, Fig.4(c) demonstrates a substantial improvement in the average number of unique topic clusters consumed per user. This finding suggests that the intent diversification strategy, by encouraging users to engage with a wider variety of novel content, leads to increased consumption diversity.
@@
(a) Novel content creator engagement.
@@
Our proposed intent diversification framework is a personalized treatment based on a personalized prediction of intent propensities for each user. To understand the effect of personalization, we conducted further analysis by comparing the treatment (with intent diversification) and control (without intent diversification) for users with different levels of exploration intent $p(Exploration)$. [^7] As the two intent probabilities sum to one, our analysis focuses only on $p(Exploration)$ here; The findings for the familiarity intent are symmetric. In particular, we looked at the following metrics: (1) Novel impressions: Number of impressions on the landing page that are from novel content creators (channels), with whom the user has not previously engaged. [^8] Note that the novelty of content creators is a personalized notion. A novel content creator for user A may not be a novel content creator for user B. This measures the overall novelty of the recommendations. (2) Novel consumptions: Number of consumed items on the landing page that are from novel content creators. (3) Novel click-through rate (CTR): Click-through-rates of the novel recommended items. This measures the relevance of the novel recommendations. Figure 5(a) - 5(c) shows the differences between treatment and control in these metrics. We observe that for pages with higher exploration intent (and consequently lower familiarity intent), the treatment (our intent diversification framework) is recommending more novel items (Fig.5(a)) than control, and the users are more likely to engaging with those novel contents (Fig.5(b) and Fig.5(c)).
All three metrics in Fig.5 exhibit a consistent monotonic trend. This suggests that our intent diversification framework improves personalization with respect to users’ intents. Specifically, for pages with a higher propensity towards exploration, our framework not only recommends an increasing proportion of high-quality novel items but also observes increased user engagement with these recommendations, as evidenced by the difference in CTR (Fig.5(c)). For example, at about 80% percentile of $p(Exploration)$, our proposed framework is showing about the same amount of novel recommendations (Fig.5(a)), but results in 4% increase in consumptions from novel contents (Fig.5(b)). This suggest that our framework is more efficient in recommending contents that are both novel and relevant to the users who are in exploration mode. Conversely, for users with lower exploration intent, the recommendation of such novel items decreases and the users are less likely to engage with them. In other words, our intent diversification framework is able to further personalize the recommendation results to align them with varying levels of user exploration intents. In Appendix A.2, we also discuss comparisons between our framework and additional baselines on exploration.
@@
4.4. Understanding User Intents
4.4.1. Model performance and feature correlation analysis
As the probabilities of exploration intent and familiarity intent sum to one, we only present the analysis for exploration intent here; the results for familiarity intent are symmetric. The user intent model has an AUC of 0.73 and is well calibrated with an average prediction to label ratio of 0.97 (Fig.6). To understand what signals are captured by the user intent model, we conducted analysis to study how each feature is correlated with the $p(Exploration)$ prediction and rank by the order of the absolute value of the correlations. Table 1 summarizes the features that have the highest absolute correlations with the predicted user exploration intent. We see that features such as “current session length” and “number of consumptions in the current session” are positively correlated with $p(Exploration)$. This is interesting and expected, as it suggests that users are more likely to explore when they have caught up existing interests (e.g. finishing consuming the newest episode of a drama series) in the current session. Features such as “average completion ratio of past consumed items” and “average total length of past consumed items” are negatively correlated with $p(Exploration)$, indicating that users who are less patient with each consumption (i.e. more likely to consume shorter items or quit before finishing the consumption) are more likely to explore. These findings are also aligned with the qualitative insights from user interviews conducted at the company.
@@ 4.4.1. Model performance and feature correlation analysis
Number of content creators interacted with
@@
Table 1. Feature analysis of user intent model.
4.4.2. User intents on different time horizons
To understand how the predicted user intents evolves over different sessions and varies across different users, we visualized one of the intent model’s predictions $p(Exploration)$ across different time horizons for a random sample of the users. Figure 7(a) shows the page-level $p(Exploration)$ over different hours of day for a random sample of 10 users. We observe that users tend to exhibit varying levels of exploration intent throughout different hours or sessions of the day, even for the same individual. This is expected, as a user’s mood and context may fluctuate between, for instance, the morning rush hours and evening family time, resulting in differing propensities to explore. Figure 7(a) shows that our intent model is able captures these variations.
Even more intriguingly, when aggregating data at the daily level rather than hourly, we observe significantly smaller fluctuations across different days, as shown in Fig.7(b). This implies that each user maintains a relatively stable baseline level of exploration intent, which may vary among individuals. While some users are inherently more likely to explore than others, within the same user, their propensity to explore remains consistent across different days. This stable nature of user-level intent provides grounds in using supervised machine learning models to predict future intents based on the user’s past behaviors.
(a) Page-level exploration intent: $p(Exploration)$ over different hours of day.
(b) User-level exploration intent: Average $p(Exploration)$ over different days.
Figure 7. User intents on different time horizons.
@@ 4.5. Effect of the Strength of Intent Diversification
Although diversity is a desired property for recommendation results, it has been established that diversity usually hurts the overall relevance of the set of recommended items (Chen et al., 2021). Therefore, the goal of the recommender system is to achieve the optimal trade-off between diversity and relevance in order to optimize long-term user experience. In our intent diversification framework, the strength of diversification is controlled by the hyperparameter $\gamma$ in Eq.(8). To understand the effect of the different strengths of intent diversification, we conduct additional experiments with varying values of $\gamma$, and measure the following four metrics: (1) Diversity, measured by number of unique clusters consumed per user; (2) Novelty, measured by novel content creator engagement; (3) Relevance, which is the average of the relevance score (predicted by a separate ML model) of all recommended items on the landing page; (4) Long-Term Metric (DAU), which is the Daily Active User (DAU) metric.
@@ 5. Conclusion
In this work, we propose a framework to diversify recommendation results based on user intents. By adapting an algorithm originally designed for search results diversification, we propose to explicitly incorporate the user’s propensities to different intents in the final stage of a recommender system, to ensure that different intents are represented in the recommendations. Through live experiments on YouTube, the world’s largest video recommendation platform, we demonstrate the effectiveness of our proposed framework in assisting users towards their intents and improving their long-term experience on the platform. In particular, we see a significant improvement in long-term metrics such as overall user enjoyment and Daily Active Users (DAU).
Our proposed framework works with any number of intents and can be applied either as a standalone diversification layer or combined with existing whole-page optimization algorithms. Although our findings is only from one platform, our framework is general and readily applicable to other recommendation platforms. In fact, our approach is more beneficial when there are not a huge amount of training data available. This is because compared to a black-box diversification algorithm, our intent-based diversification framework is able to better capture the underlying data generation process (i.e. an intent-driven decision-making process). It improves learning efficiency given the same amount of training data as a result. Therefore, smaller platforms should benefit even more from our framework. In addition, the intent prediction model is much smaller in size compared with the other models used by the platform. Therefore, it introduces minimal engineering cost in implementing our framework.
@@
In the live experiments, we selected the intents based on business considerations by the platform. A future research direction is to automatically learn a taxonomy of user intents that can directly serve as input to the intent diversification framework toward optimizing long-term user experience on the platform.
Appendix A Additional experiment results
A.1. Experiments on additional intents
We define a creator intent as the user’s propensity to engage with content from a specific creator on the platform (i.e., a channel on YouTube). Considering the fact that YouTube hosts hundreds of millions of creators, the set of intents $\mathcal{V}$ is extremely large in this case. We were able to conduct A/B testing on the intent diversification framework with the creator intents.
Figure 9 below summarizes the results. There is a significant 0.05% increase in Daily Active Users (DAU) from incorporating creator intents (Fig.9(a)). This is accompanied by a 0.04% increase in site-wide consumption count (Fig.9(b)). In terms of intent-related metrics, we see that there is a significant 0.56% increase in novel creator consumption, 0.10% increase in unique user-creator pairs, 0.13% increase in unique user-cluster pairs, and 0.30% repeated novel consumption, suggesting that our creator-intent-aware treatment facilitates the discovery and engagement with novel content creators. As a result, we also see an increase in consumption diversity which is measured by unique user-cluster pairs.
(a) Daily Active Users.
(b) Site-wide consumption count.
(c) Novel creator consumption.
(d) Unique user-creator pairs.
(e) Unique user-cluster pairs.
(f) Repeated novel consumption.
Figure 9. Creator-intent related metrics.
We also observed positive A/B experiment results when modeling user intents around the visit length on the platform within our intent diversification framework. Specifically, we build an ML model to predict the probability that a user would prefer to watch a long versus a short video on the current recommendation page. By incorporating these visit length intents, we achieved a significant increase of +0.06% in overall user enjoyment and a notable +0.59% increase in landing page consumption, as shown in Fig.10. These outcomes confirm our framework’s ability to effectively generalize and scale across a broad set of user intents.
(a) Overall user enjoyment.
(b) Landing page consumption.
Figure 10. Visit length intent related metrics.
A.2. Additional baselines on exploration
Our intent diversification framework with exploration and familiarity intents also outperforms existing state-of-the-art baselines for exploration in recommender systems. The existing production recommender system at the platform (which serves as control in our experiments) has already incorporated several state-of-the-art approaches for boosting novelty and diversity in the recommendation results. For diversification, the control group has incorporated a treatment similar to Wilhelm et al. (2018) and Wang et al. (2022); For novelty, the control group has incorporated a treatment similar to Chen et al. (2021). For general diversification algorithms, the platform has experimented with approaches ranging from rule-based heuristics to model-based diversification (Wilhelm et al., 2018). Our proposed intent diversification algorithm is able to provide further improvements (in DAU etc.) on top of these existing state-of-the-art approaches for diversified and novel recommendations.